Find the y-intercept of the line with the equation 2x-6y=15
The answers for me to pick from are -5/2, 15/2, -15/2, 15
I came up with the answer -5/2, here is my work...
15=2x-6y
15 divided by -6 = -5/2
Answer -5/2
Am I correct?

Answer by stanbon(75887)

Find the y-intercept of the line with the equation 2x-6y=15
----------
Here are two ways to work the problem:
1st: To find the y-intercept let x=0:
You get -6y = 15; y = -5/2
-------------------
2nd: Put the equation in slope-intercept form:
6y = 2x-15
Divide both sides by 6 to get:
y = (1/3)x-(5/2)
y-intercept is -5/2
===============
Cheers,
Stan H.